A Complete Bibliography of Publications in Applied Mathematics and Computation: 2025–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 September 2024 Version 1.01

Title word cross-reference

1 [JSZ25]. 3 [MR25]. 4 [LZ25].

-factorization [JSZ25]. -independence [LZ25].

1 [Ano25a]. 15 [Ano25b].

2025 [Ano25a, Ano25b].

Acoustics [ALC^{+}25]. action [MMNP25]. actuator [SNP25]. adaptive [ZLA25, ZLCL25]. advection [PP25]. advection-diffusion [PP25]. against [SNP25]. algorithms [LTH^{+}25]. Allen [OGPFVV25]. allocation [RS25]. Analysis [FBK25, ALC^{+}25]. applied [TNL25]. approach [DAG25]. associative [SNP25]. asynchronous [MCH25]. Attack [MCH25]. Attack-compensated [MCH25]. attacks [ZLCL25]. attitude [ZLCL25]. avoidance [ZLA25].

B [PP25]. B-splines [PP25]. background [RS25]. Barrier [ZLA25]. Barrier-function [ZLA25]. based [PP25, ZLA25]. basis [TNL25]. bidirectional [SNP25]. blind [IAK25]. Board [Ano25c, Ano25d]. boundary [ALC^{+}25, OGPFVV25].

Cahn [OGPFVV25]. collision [ZLA25]. collocation [PP25]. compact [MR25]. compensated [MCH25]. competitive [FBK25]. complete [JSZ25, LYM25]. conditions [ALC^{+}25, OGPFVV25]. consensus [GPL25]. conservation [TNL25]. constrained [GPL25]. control [GPL25, MCH25, SNP25, ZLA25, ZLCL25]. cross [FBK25]. cross-immunity [FBK25].

D [MR25]. deblurring [IAK25]. deception [ZLCL25]. delayed [SZY25, SNP25]. description [MMNP25]. deterministic [DAG25]. difference [OGPFVV25]. differential [GPL25]. diffusion [PP25]. Dirichlet [ALC^{+}25]. disease [FBK25]. distribution [MP25]. driven [GPL25]. Dynamic [GPL25, ZLCL25]. dynamics [Che25].

Editorial [Ano25c, Ano25d]. effects [FBK25]. eigenvalue [LYM25, MP25]. Epidemic [FBK25]. epidemics [MMNP25]. equation [LTH^{+}25, PP25]. equations [ALC^{+}25, MR25, OGPFVV25]. Erdos [MMNP25]. estimates [TNL25]. estimation [SZY25]. event [GPL25, ZLCL25]. event-driven [GPL25]. event-triggered [ZLCL25]. Exploring [MMNP25]. exponential [MR25]. extrema [LZ25].

factorization [JSZ25]. failure [SNP25]. feedback [MCH25, SNP25]. field [MMNP25]. finite [OGPFVV25]. finite-difference [OGPFVV25]. form [MR25]. fourth [PP25]. fragile [SNP25]. framework [MAMFBM^{+}25]. frictions [RS25]. function [TNL25, ZLA25].

games [GPL25]. generalized [LZ25]. given [LZ25]. graphical [GPL25]. graphs [LYM25, LZ25, MP25]. growth [BDL25].

high [MR25, ZLA25]. high-order [MR25, ZLA25]. hyperbolic [PP25]. hypergraph [JSZ25]. hypergraphs [MAMFBM^{+}25].

image [IAK25]. immunity [FBK25].

Impenetrability [ALC^{+}25]. implicit [MR25]. independence [LZ25]. influence [DAG25]. insurable [RS25]. investment [RS25].

January [Ano25a, Ano25b].

Laplacian [LYM25, MP25]. law

[LTH^{+}25, MMNP25]. laws [TNL25]. least [LYM25]. Level [IAK25, MR25]. limits [MMNP25]. linear [MR25, TNL25]. Local [LTH^{+}25]. Lotka [Che25].

mass [MMNP25]. mean [MMNP25]. mechanism [Che25]. memory [SNP25]. memristive [SZY25, SNP25]. metastatic [BDL25]. method [IAK25, PP25]. methods [TNL25]. mimetic [OGPFVV25]. Minimizing [LYM25]. model [Che25]. Modelling [DAG25]. models [BDL25]. multiagent [GPL25].

networks [MMNP25, SZY25, SNP25].

neural [SZY25, SNP25]. Non

[SNP25, JSZ25]. Non-fragile [SNP25]. non-uniform [JSZ25]. nonlinear [ALC^{+}25, LTH^{+}25, RS25, ZLA25]. nonlinearity [LTH^{+}25]. nonperiodic [OGPFVV25]. note [JSZ25]. number [LZ25]. Numerical [ALC^{+}25, BDL25].

operators [OGPFVV25]. Optimal

[RS25, GPL25]. order

[LZ25, MR25, PP25, SZY25, ZLA25].

output [MCH25, SNP25]. output-feedback [SNP25].

parabolic [MR25]. Pattern [Che25]. periodic [OGPFVV25]. polynomial [PP25]. portfolio [RS25]. poverty [DAG25]. power [LTH^{+}25]. predefined [ZLCL25]. predefined-time [ZLCL25]. preserving [LTH^{+}25]. prevalence [DAG25]. probability [MCH25]. problems [IAK25]. prostitution [DAG25].

quarantine [FBK25]. quasi [MR25]. quasi-linear [MR25]. QUAV [ZLCL25].

radial [TNL25]. Reachable [SZY25]. Redefined [PP25]. Rényi [MMNP25]. respiratory [FBK25]. risk [RS25].

scalar [TNL25]. scale [MAMFBM^{+}25]. scheme [MR25]. Schrödinger [LTH^{+}25]. second [SZY25]. second-order [SZY25]. set [SZY25]. signed [LYM25]. sojourn [MCH25]. solution [BDL25]. Solving [OGPFVV25, PP25]. spectral [LZ25]. splines [PP25]. Stability [TNL25]. state [GPL25]. state-constrained [GPL25]. stochastic [DAG25, MCH25]. structure [LTH^{+}25]. structure-preserving [LTH^{+}25]. sum [GPL25]. switching [MCH25]. system [FBK25]. systems [MCH25, ZLA25].

taxis [Che25]. thresholds [FBK25]. time [ZLCL25]. Topological [MAMFBM^{+}25]. tracking [ZLA25]. trajectory [ZLA25]. treatment [BDL25]. triggered [ZLCL25]. tumor [BDL25]. Two [IAK25, MR25]. Two-Level [IAK25, MR25].

unicyclic [MP25]. uniform [JSZ25, PP25]. unknown [ZLCL25]. using [OGPFVV25].

Volterra [Che25]. vs [DAG25].

wave [ALC^{+}25].

zero [GPL25]. zero-sum [GPL25].

References

Alcantara:2025:NAN

[ALC^{+}25] Adriano A. Alcântara, Juan B.

Límaco, Bruno A. Carmo, Ronald R. Guardia, and Mauro A. Rincon. Numerical analysis for nonlinear wave equations with boundary conditions: Dirichlet, acoustics and impenetrability. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004703.

Anonymous:2025:Ja

[Ano25a]  Anonymous.  1 January

2025. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:Jb

[Ano25b] Anonymous. 15 January 2025. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Anonymous:2025:EBa

[Ano25c]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005101.

Anonymous:2025:EBb

[Ano25d]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324005241.

Bulai:2025:NSM

[BDL25] I. M. Bulai, M. C. De Bonis, and C. Laurita. Numerical solution of metastatic tumor growth models with treatment. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004491.

Chen:2025:PDL

[Che25] Mengxin Chen. Pattern dynamics of a Lotka–Volterra model with taxis mechanism. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004788.

Divya:2025:MPP

[DAG25] G. Divya, S. Athithan, and Mini Ghosh. Modelling the prevalence of prostitution under the influence of poverty: a deterministic vs. stochastic approach. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004910.

Fome:2025:ACR

[FBK25] Anna Daniel Fome, Wolfgang Bock, and Axel Klar. Analysis of a competitive respiratory disease system with quarantine: Epidemic thresholds and crossimmunity effects. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004296.

Guo:2025:DED

[GPL25] Siyu Guo, Yingnan Pan, and Hongyi Li. Dynamic event-driven optimal consensus control for stateconstrained multiagent zerosum differential graphical games. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004405.

Iqbal:2025:TLM

[IAK25] Azhar Iqbal, Shahbaz Ahmad, and Junseok Kim. Two-level method for blind image deblurring problems. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004697.

Jiang:2025:NFN

[JSZ25] Taijiang Jiang, Qiang Sun, and Chao Zhang. A note on the 1-factorization of nonuniform complete hypergraph. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004685.

Luo:2025:LSP

[LTH^{+}25] Fangwen Luo, Qiong Tang, Yiting Huang, Yanhui Ding, and Sijia Tang. Local structure-preserving algorithms for the nonlinear Schrödinger equation with power law nonlinearity. Applied Mathematics and Computation, 484(??): ??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004478.

Li:2025:MLL

[LYM25] Dan Li, Minghui Yan, and Jixiang Meng. Minimizing the least Laplacian eigenvalue of signed complete graphs. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004636.

Li:2025:SEG

[LZ25] Shuchao Li and Zihan Zhou. On spectral extrema of graphs with given order and generalized 4-independence number. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400479X.

Molina-Abril:2025:TSF

[MAMFBM^{+}25] H. Molina-Abril, M. J. Morón-Fernández, M. Benito-Marimón, F. Díaz del Río, and P. Real. Topological scale framework for hypergraphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004508.

Mei:2025:ACA

[MCH25] Xu Mei, Jun Cheng, and Wentao Huang. Attackcompensated asynchronous output feedback control for stochastic switching systems with sojourn probability. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004855.

Munoz:2025:ELL

[MMNP25] Francisco J. Muñoz, Luca Meacci, Juan Carlos Nuño, and Mario Primicerio. Exploring the limits of the law of mass action in the mean field description of epidemics on Erdős–Rényi networks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004806.

Moon:2025:LED

[MP25] Sunyo Moon and Seungkook Park. Laplacian eigenvalue distribution for unicyclic graphs. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004831.

Mittal:2025:TLI

[MR25] Kajal Mittal and Rajendra K. Ray. Two-level implicit high-order compact scheme in exponential form for 3D quasilinear parabolic equations. Applied Mathematics and Computation, 485(??):??, January 15, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004454.

Orizaga:2025:SAC

[OGPFVV25] Saulo Orizaga, Gilberto González-Parra, Logan Forman, and Jesus Villegas-Villanueva. Solving Allen– Cahn equations with periodic and nonperiodic boundary conditions using mimetic finite-difference operators. Applied Mathematics and Computation, 484(??):??, January 1, 2025. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004545.

Palav:2025:RFO

[PP25] Mansi S. Palav and Vikas H.

Pradhan. Redefined fourth order uniform hyperbolic polynomial B-splines based collocation method for solving advection-diffusion equation. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004533.

Ramirez:2025:OII

[RS25] Hugo E. Ramírez and Rafael Serrano. Optimal investment with insurable background risk and nonlinear portfolio allocation frictions. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004843.

Suvetha:2025:NFO

[SNP25] R. Suvetha, J. J. Nieto, and P. Prakash. Nonfragile output-feedback control for delayed memristive bidirectional associative memory neural networks against actuator failure. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S009630032400482X.

Shen:2025:RSE

[SZY25] Yi Shen, Jiemei Zhao, and Liqi Yu. Reachable set estimation of delayed secondorder memristive neural networks. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004557.

Tominec:2025:SER

[TNL25] Igor Tominec, Murtazo Nazarov, and Elisabeth Larsson. Stability estimates for radial basis function methods applied to linear scalar conservation laws. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0096300324004818.

Zhang:2025:BFB

[ZLA25] Lili Zhang, Chenglong Liu, and Liwei An. Barrierfunction based adaptive trajectory tracking control for high-order nonlinear systems with collision avoidance. Applied Mathematics and Computation, 484 (??):??, January 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S009630032400465X.

Zhu:2025:DET

[ZLCL25] Guanchao Zhu, Min Luo, Guozeng Cui, and Ze Li. Dynamic event-triggered predefined-time adaptive attitude control for a QUAV with unknown deception attacks. Applied Mathematics and Computation, 485 (??):??, January 15, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0096300324004648.